OpenStudy (anonymous):

Expand each product (x-9)(x+6)

OpenStudy (anonymous):

Use FOIL, First, Outter, Inner,Last. (X-9)(x+6)... First is x*x =x^2. Outter is x*6=6x. Inner is -9*x=-9x. Last is -9*6=-54 Your expansion is x^2+6x-9x-54... Don't forget to combine like terms (the 6x and-9x)... x^2-3x-54 is your answer